solve each system of equations by substitution method 8x+3y=-7
x=3y+15

Answer by checkley79(3341)

8x+3y=-7
x=3y+15
8(3Y+15)+3Y=-7
24Y+120+3Y=-7
27Y=-7-120
27Y=-127
Y=-127/27
Y=-4.7037 ANS.
X=3*-4.7037+15
X=-14.111+15
X=.8889 ANS.
PROOF:
8*.8889+3*-4.7037=-7
7.111-14.111=-7
-7=-7